# Supervised Graph Contrastive Learning for Gene Regulatory Networks

### v4

@@ C.1 Details on the TCGA Dataset
In this study, for gene expression data, we used normalized count data from the TCGA TARGET GTEx study provided by UCSC Xena (UCSC Xena (2016)) for the TCGA dataset. For the LINCS dataset, we used normalized gene expression data from the LINCS L1000 GEO dataset (GSE92742) (Subramanian (2017)).
@@ C.3 Extraction of Gene Label Data for BP/CC Tasks using Gene Ontology
In this study, for Biological Process (BP) classification and Cellular Component (CC) classification in downstream tasks, we performed multi-label annotation for each of the 975 genes constituting the GRN, based on terms obtained from the GO database. For BP labels, three categories whose importance is known were used: ’Metabolism’, ’Signal Transduction’, and ’Cellular Organization’  (Paolacci et al. (2009)). Similarly, for CC labels, four categories were used: ’Nucleus’, ’Mitochondrion’, ’Endoplasmic Reticulum’, and ’Plasma Membrane’  (Costa et al. (2010)). These categories were selected to cover the major functions of the GRN.
@@
    Batch retrieve terms associated with each gene using the MyGene.info API (Xin et al. (2015)) and the OBO file (available at https://geneontology.org/docs/download-ontology/).
@@
    Aggregate multi-labels for the target genes using the GOATOOLS library  (Ashburner et al. (2000)).
@@ D.1 Taxonomy of Gene Regulatory Network Estimation
Estimating accurate gene regulatory networks (GRNs) is crucial for elucidating cellular processes and disease mechanisms. Methods for computationally estimating GRNs from gene expression data can be categorized into correlation-based (Langfelder & Horvath (2008)), mutual information-based  (Margolin et al. (2006)), probabilistic graphical models-based such as Bayesian networks  (Friedman et al. (2000)), and deep learning-based approaches  (Shu et al. (2021)). The number of experimentally validated GRNs is limited. Therefore, GRN estimation methods need the ability to account for measurement errors and appropriately capture non-linear and multimodal interactions between genes while mitigating overfitting. Thus, we adopted a method combining Bayesian network estimation using multiple sampling and non-parametric regression (Imoto et al., 2002) to identify patient- or sample-specific GRNs (Nakazawa et al., 2021).
@@ D.2 Details of Estimating GRNs
The Bayesian network is estimated by learning the relationships with parent nodes based on the model described above and the parameters of the conditional distributions. In this study, the score function used for searching the structure of the Bayesian Network can be analytically derived using Laplace approximation (Imoto et al., 2002). Since the problem of finding a Directed Acyclic Graph (DAG) that maximizes this score is NP-hard, we performed structure search using a heuristic structure estimation algorithm, the greedy hill-climbing (HC) algorithm  (Imoto et al. (2003)). Furthermore, to ensure the reliability of the estimation results by the HC algorithm, we performed multiple sampling runs.
@@
Furthermore, the feature of the $i$-th node $X_{i}^{\mathcal{V}}$ is the gene expression level of each sample, and the feature of an edge from $i$ to $j$ (designated as the $k$-th edge) is the realization of the learned B-spline curve $X_{k}^{\mathcal{E}}=m_{ij}(x_{j})$. Patient-specific networks using such edge features have led to the discovery of subtypes that correlate more strongly with prognosis than existing subtypes (Nakazawa et al., 2021). Additionally, using differences in edge features between patients to extract patient-specific networks has been reported to contribute to the identification of novel diagnostic and therapeutic marker candidates in diseases such as idiopathic pulmonary fibrosis  (Tomoto et al. (2024)) and chronic nonbacterial osteomyelitis  (Yahara et al. (2022))
For GRN estimation, we used INGOR (version 0.19.0), a software that estimates Bayesian Networks based on B-spline regression, and executed it on the supercomputer Fugaku. INGOR is based on SiGN-BN  (Tamada et al. (2011)), a software that similarly estimates Bayesian Networks using B-spline regression, and achieves faster estimation by optimizing parallel computation on Fugaku. In all GRN estimations, the number of sampling runs was set to 1000, and the threshold for adopting edges was set to 0.05. Since network estimation with a large amount of sample data can lead to Out of Memory errors, the upper limit of gene expression data used for network estimation was set to 3000 (especially for LINCS data). All other hyperparameters related to network estimation used the default settings of SiGN-BN. The number of Fugaku nodes and the required execution times are shown in Table 7. In addition, the number of parallel threads was set to 4 for estimations using TCGA data and 2 for LINCS data.
@@ E.2 Hyperparameters Settings
In this study, we used Optuna (Akiba et al., 2019), a Bayesian optimization tool, for hyperparameter search in pre-training. The search space for all models included the AdamW learning rate $\mathrm{lr}\in[10^{-5},10^{-3}]$, batch size $\mathrm{batch\_size}\in\{4,\,8\}$, and model-specific hyperparameters. Model-specific hyperparameters were the temperature parameter $\tau\in\{0.25,\,0.5,\,0.75,\,1.0\}$ for GraphCL, GRACE, and SupGCL, and the global-hop parameter $k\in\{1,2,3\}$ for SGRL. The graph embedding dimension was unified to 64 for all models. See Appendix F for a discussion on embedding dimension. For training the pre-trained models, the data was split into training and validation sets at an $8:2$ ratio, and the validation loss was used as the metric for various decisions. The optimal hyperparameters determined by actual hyperparameter tuning are shown in Table 10.
@@
E.3 Computational Environment and Computation Time
All experiments were conducted on an NVIDIA H100 SXM5 (95.83 GiB), and the computation time for each model is shown in Table 11. Please note that although the number of training steps is based on 3000 epochs, the actual training time varies due to early stopping using the validation data.
Table 11: Pre-training computation time for each cancer type

### v5

@@ C.1 Details on the TCGA Dataset
In this study, for gene expression data, we used normalized count data from the TCGA TARGET GTEx study provided by UCSC Xena (UCSC Xena, 2016) for the TCGA dataset. For the LINCS dataset, we used normalized gene expression data from the LINCS L1000 GEO dataset (GSE92742) (Subramanian, 2017).
@@ C.2.1 Preparing LINCS Dataset
C.2.2 Discussion for the Quality of LINCS Dataset
It is well-documented that LINCS datasets contain significant noise, particularly off-target effects (Subramanian et al., 2017b). While aggregation methods such as Consensus Gene Signatures (CGS) have been proposed to isolate consistent on-target effects (Subramanian et al., 2017b), they inherently reduce the effective sample size—a trade-off we sought to avoid in this study. Despite the use of instance-level data without such aggregation, SupGCL demonstrated superior predictive performance on downstream tasks, especially graph-level classification, compared to existing methods. Furthermore, our enrichment analysis(Appendix  J) reveals that SupGCL successfully acquired latent representations with high biological interpretability. These findings suggest that SupGCL is capable of robust pretraining even when learning from noisy teacher data.
@@ C.3 Extraction of Gene Label Data for BP/CC Tasks using Gene Ontology
In this study, for Biological Process (BP) classification and Cellular Component (CC) classification in downstream tasks, we performed multi-label annotation for each of the 975 genes constituting the GRN, based on terms obtained from the GO database. For BP labels, three categories whose importance is known were used: ’Metabolism’, ’Signal Transduction’, and ’Cellular Organization’  (Paolacci et al., 2009). Similarly, for CC labels, four categories were used: ’Nucleus’, ’Mitochondrion’, ’Endoplasmic Reticulum’, and ’Plasma Membrane’  (Costa et al., 2010). These categories were selected to cover the major functions of the GRN.
@@
    Batch retrieve terms associated with each gene using the MyGene.info API (Xin et al., 2015) and the OBO file (available at https://geneontology.org/docs/download-ontology/).
@@
    Aggregate multi-labels for the target genes using the GOATOOLS library  (Ashburner et al., 2000).
@@ D.1 Taxonomy of Gene Regulatory Network Estimation
Estimating accurate gene regulatory networks (GRNs) is crucial for elucidating cellular processes and disease mechanisms. Methods for computationally estimating GRNs from gene expression data can be categorized into correlation-based (Langfelder & Horvath, 2008), mutual information-based  (Margolin et al., 2006), probabilistic graphical models-based such as Bayesian networks  (Friedman et al., 2000), and deep learning-based approaches  (Shu et al., 2021). The number of experimentally validated GRNs is limited. Therefore, GRN estimation methods need the ability to account for measurement errors and appropriately capture non-linear and multimodal interactions between genes while mitigating overfitting. Thus, we adopted a method combining Bayesian network estimation using multiple sampling and non-parametric regression (Imoto et al., 2002) to identify patient- or sample-specific GRNs (Nakazawa et al., 2021).
@@ D.2 Details of Estimating GRNs
The Bayesian network is estimated by learning the relationships with parent nodes based on the model described above and the parameters of the conditional distributions. In this study, the score function used for searching the structure of the Bayesian Network can be analytically derived using Laplace approximation (Imoto et al., 2002). Since the problem of finding a Directed Acyclic Graph (DAG) that maximizes this score is NP-hard, we performed structure search using a heuristic structure estimation algorithm, the greedy hill-climbing (HC) algorithm  (Imoto et al., 2003). Furthermore, to ensure the reliability of the estimation results by the HC algorithm, we performed multiple sampling runs.
@@
Furthermore, the feature of the $i$-th node $X_{i}^{\mathcal{V}}$ is the gene expression level of each sample, and the feature of an edge from $i$ to $j$ (designated as the $k$-th edge) is the realization of the learned B-spline curve $X_{k}^{\mathcal{E}}=m_{ij}(x_{j})$. Patient-specific networks using such edge features have led to the discovery of subtypes that correlate more strongly with prognosis than existing subtypes (Nakazawa et al., 2021). Additionally, using differences in edge features between patients to extract patient-specific networks has been reported to contribute to the identification of novel diagnostic and therapeutic marker candidates in diseases such as idiopathic pulmonary fibrosis  (Tomoto et al., 2024) and chronic nonbacterial osteomyelitis  (Yahara et al., 2022)
For GRN estimation, we used INGOR (version 0.19.0), a software that estimates Bayesian Networks based on B-spline regression, and executed it on the supercomputer Fugaku. INGOR is based on SiGN-BN  (Tamada et al., 2011), a software that similarly estimates Bayesian Networks using B-spline regression, and achieves faster estimation by optimizing parallel computation on Fugaku. In all GRN estimations, the number of sampling runs was set to 1000, and the threshold for adopting edges was set to 0.05. Since network estimation with a large amount of sample data can lead to Out of Memory errors, the upper limit of gene expression data used for network estimation was set to 3000 (especially for LINCS data). All other hyperparameters related to network estimation used the default settings of SiGN-BN. The number of Fugaku nodes and the required execution times are shown in Table 5. In addition, the number of parallel threads was set to 4 for estimations using TCGA data and 2 for LINCS data.
@@ E.2 Hyperparameters Settings
In this study, we used Optuna (Akiba et al., 2019), a Bayesian optimization tool, for hyperparameter search in pre-training. The search space for all models included the AdamW learning rate $\mathrm{lr}\in[10^{-5},10^{-3}]$, batch size $\mathrm{batch\_size}\in\{4,\,8\}$, and model-specific hyperparameters. Model-specific hyperparameters were the temperature parameter $\tau\in\{0.25,\,0.5,\,0.75,\,1.0\}$ for GraphCL, GRACE, and SupGCL, and the global-hop parameter $k\in\{1,2,3\}$ for SGRL. The graph embedding dimension was unified to 64 for all models. See Appendix G for a discussion on embedding dimension. For training the pre-trained models, the data was split into training and validation sets at an $8:2$ ratio, and the validation loss was used as the metric for various decisions. The optimal hyperparameters determined by actual hyperparameter tuning are shown in Table 8.
@@
E.3 Fine-tuning Settings
This section described the fine-tuning setting of downstream-task. The overview is in Table 9.
Table 9: Description of the downstream tasks
Task Task Type Metrics
Node-Level Task
[Rel.]: Cancer relation Classification (binary) Accuracy
Graph-Level Task
E.3.1 Graph-Level Task
Hazard Prediction :
$h(t\mid{\bm{x}})=h_{0}(t)\exp\bigl(\bm{\beta}^{\top}{\bm{x}}\bigr)$
Subtype Classification:
For the subtype classification task, we created a classification model using a 5-class softmax function and trained it using multi-class cross-entropy. Furthermore, performance was evaluated using Accuracy and Macro F1-score. The F1-score results are shown in Appendix G.
E.3.2 Node‐Level Task
BP./CC. Classification :
In Biological Process (BP) classification, three categories for each gene—”metabolism,” ”signal transduction,” and ”cellular organization”—are predicted as a multi-hot vector. In Cellular Component (CC) classification, four categories—”nucleus,” ”mitochondria,” ”endoplasmic reticulum,” and ”plasma membrane”—are predicted as a multi-hot vector. The model was structured using a sigmoid function for each category, and training was performed using binary cross-entropy for each respective category. Performance was evaluated using Subset Accuracy, Macro F1-score, and Jaccard Index as evaluation metrics. The Macro F1-score and Jaccard Index results are shown in Appendix G.
Cancer Rel. Classification
In cancer-related gene classification, 106 genes defined as positive by OncoKB were labeled as ”positive,” and all other genes were labeled as ”negative” for binary classification. A model was created to estimate negative and positive cases using a sigmoid function, and training was performed using binary cross-entropy. Performance was evaluated using Accuracy and F1-score as evaluation metrics. The F1-score results are shown in Appendix G.
Appendix F Empirical Runtime and Computational Complexity of SupGCL
In this section, we first report the empirical runtime of the proposed method and the comparative methods, and then discuss the computational complexity of the proposed and conventional GCL methods.
F.1 Computational Environment and Computation Time
All experiments were conducted on an NVIDIA H100 SXM5 (95.83 GiB), and the computational run time for each model is shown in Table 10. Please note that although the number of training steps is based on 3000 epochs, the actual training time varies due to early stopping using the validation data.
Table 10: Pre-training computational run time for each cancer type